2. You are interested in purchasing a Honda Acura. The salesperson indicated that the price of the car was either $26,000 cash or $6,500 at the end of each of 5 years. Compute the effective-interest rate to the nearest percent that you will pay if you choose to make the five annual payments. (Round answer to 0 decimal places, e.g. 79%) Interest Rate Answe 3. Mickey Michen will invest $6,420 a year for 24 years in a fund that will earn 5% annual interest. If the first payment into the fund occurs today, what amount will be in the fund in 24 years? (Round factor values to 5 decimal places, eg. 1.25124 and final answers to 0 decimal places, e.g. 458581.) Answer: 4. On January 1, 2017, Sheffield Corp. issued ten-year bonds with a face amount of $5800000 and a stated interest rate of 8% payable annually on January 1 The bonds were priced to yield 10%. Present value factors are as follows: At 8% 0.463 6.710 At 10% 0.386 6.145 Present value of 1 for 10 periods Present value of an ordinary annuity of I for 10 periods The total issue price of the bonds was Answer: 5. On December 31, 2017, Bramble Company acquired a computer from Plato Corporation by issuing a s592,000 zero-interest-bearing note, payable in full on December 31, 2021. Bramble Companys credit rating permits it to borrow funds frorn its several lines of credit at 12%. The computer is expected to have a 5-year life and a $66,000 salvage value.
Answer #1

2)We can use rate formuale in excel to solve this
=rate(nper,pmt,pv,fv,type)
=rate(5,-6500,26000,0,0)
=8%

3)we use fv formulae in excel to solve this
=fv(rate,nper,pmt,pv,type)
=fv(5%,24,6240,0,1)
=291577

4)bond price=(coupon * pvA(10%,10))+(face value*PV(10%,10)
=(5800000*8%*6.145)+(5800000*0.386)
=5090080